Calculate the molar solubility of CuSO. (Ksp = 8.1 x 10-16) a) in water b) in 0.10 M CUNO3(aq) solution
Answer #1

Q2. (a) Cu2SO3 (s) \rightleftharpoons 2 Cu+ (aq) + SO32- (aq)

Let molar solubility of Cu2SO3 = s

[Cu+] = 2s

[SO32-] = s

Ksp Cu2SO3 = [Cu+]2[SO32-]

8.1 x 10-16 = (2s)2 * (s)

8.1 x 10-16 = 4s3

s = (8.1 x 10-16 / 4)1/3

s = 5.9 x 10-6 M

molar solubility of Cu2SO3 = s = 5.9 x 10-6 M

(b) Initial concentration of Cu+ = 0.10 M

[Cu+] = 2s + 0.10 M

[SO32-] = s

Ksp Cu2SO3 = [Cu+]2[SO32-]

8.1 x 10-16 = (2s + 0.10 M)2 * (s)

Let s << 0.10 M

8.1 x 10-16 = (0.10 M)2 * (s)

s = (8.1 x 10-16) / (0.10 M)2

s = 8.1 x 10-14 M

molar solubility of Cu2SO3 = s = 8.1 x 10-14 M